OpenStudy (anonymous):

One pipe fills a tank in 1 hour while another pipe can fill the same tank in 2 hours. How long will it take to fill the tank if both pipes are used

OpenStudy (lgbasallote):

use the equation \[\frac{t}{1} +\frac{t}{2} = 1\]

OpenStudy (lgbasallote):

where t = combined time

OpenStudy (anonymous):

in one hour, both pipes will fill the tank's 1+1/2 part=3/2 part so, the tank can be filled in 2/3 hours=40 minutes

OpenStudy (anonymous):

my answer: one pipe fills tank in 1 hr and one pipe in 2 hrs. in 40 minutes, first pipe fills tank (2/3 full) and 2nd pipe fills tank 1/3 full. Total time = 40 minutes same as Arnab's